# A new bump in the night: Evidence of a new feature in the binary black hole mass distribution at $70~{}M_{\odot}$ from gravitational-wave observations

### v1

@@ Abstract
We analyze the confident binary black hole (BBH) detections from the third Gravitational-Wave Transient Catalog (GWTC-3) with an alternative mass population model in order to capture features in the mass distribution beyond the Powerlaw + Peak model. We find that the peak of a second power law characterizes the $\sim 30-35~{}M_{\odot}$ bump, such that the data marginally prefers a mixture of two power laws for the mass distribution of binary components over a Powerlaw + Peak model with a Bayes Factor $\log_{10}\mathcal{B}$ of 0.1. This result may imply that the $\sim 30-35~{}M_{\odot}$ feature represents the onset of a second population of BBH mergers (e.g. from a dynamical formation channel) rather than a specific mass feature over a broader distribution. When an additional Gaussian bump is allowed within our power law mixture model, we find a new feature in the BH mass spectrum at $\sim 65-70~{}M_{\odot}$. This new feature may be consistent with hierarchical mergers, and constitute $\sim 2\%$ of the BBH population. This model also recovers a maximum mass of $58^{+30}_{-14}~{}M_{\odot}$ for the second power law, consistent with the onset of a pair-instability supernova mass gap.
gravitational waves, populations
1 Introduction
Since the first detection of a binary black hole (BBH) merger in gravitational waves (GW; Abbott et al. 2016a), the field of GW astronomy has entered the era of population studies. During the course of the first three observing runs conducted by LIGO and Virgo, 69 high significance BBH mergers have been detected (Abbott et al., 2023a), and hundreds are expected to be identified throughout the ongoing fourth LIGO-Virgo-KAGRA (LVK) observing run.
Multiple formation channels have been proposed to explain the observed population of compact objects beyond the standard isolated binaries scenario, including formation pathways requiring dynamical interactions in dense star clusters (e.g. Rodriguez et al. 2015, 2019), in the disks of Active Galactic Nuclei (AGN; e.g. McKernan et al. 2012; Antonini & Rasio 2016; McKernan et al. 2018) or in the center of dwarf galaxies (Conselice et al., 2020; Palmese & Conselice, 2021), black holes from metal-free Population III stars (Pop III; e.g. Kinugawa et al. 2014; Hartwig et al. 2016), and primordial black holes (e.g. Carr & Hawking 1974; Clesse & García-Bellido 2017; Tsai et al. 2021); see Mapelli (2021); Mandel & Farmer (2022) for reviews.
@@ 1 Introduction
Multiple options for fitting the BBH mass distribution (and more broadly other population properties) have been proposed. The second and third GWTC fiducial analyses include a broken power law or a power law plus peak (Abbott et al., 2021a, b). Recent studies have also modeled the distribution with physics-motivated models (Baxter et al., 2021) or non-parametric approaches (Mandel et al., 2016; Abbott et al., 2021b; Ray et al., 2024; Magaña Hernandez & Ray, 2024; Farah et al., 2024a).
Motivated by the evidence for multiple BH populations, the presence of a peak at $\sim 30-35~{}M_{\odot}$ in the primary mass distribution, which is seemingly too narrow or at a location which is unrealistic to be explained by various formation channels (e.g. Golomb et al. 2023), the prediction for BBH from Pop III stars to peak around these values (Kinugawa et al., 2014), and the possibility of those being the seed of galaxies’ central black holes still residing in ultra-dwarf galaxies in the nearby Universe (Conselice et al., 2020; Palmese & Conselice, 2021), we explore whether the $\sim 30-35~{}M_{\odot}$ excess can be explained by a second power law rather than a peak over a broader power law, using the third GW Transient Catalog (GWTC-3; Abbott et al. 2023a).
Evidence for multiple BH populations includes the presence of a peak in the primary mass distribution at $\sim 30-35~{}M_{\odot}$, which is seemingly too narrow or too unrealistically located to to be explained by various formation channels Golomb et al. 2023). Furthermore, BBH from Pop III stars, which may be the seeds of of galaxies’ central black holes still residing in ultra-dwarf galaxies in the nearby Universe, are also predicted to peak around these values (Kinugawa et al., 2014; Conselice et al., 2020; Palmese & Conselice, 2021). Motivated by these theories, we explore whether the $\sim 30-35~{}M_{\odot}$ excess can be explained by a second power law rather than a peak over a broader power law by using the third GW Transient Catalog (GWTC-3) Abbott et al. 2023a).
Given that we are considering this possibility for a formation channel which is possibly dynamical, we also explore the presence of an additional high mass hierarchical population born from the peak of the power law.
We present our method in §2, show our results for GWTC-3 in §3, present validity tests using simulated GW events in §4, and discuss spin measurements in §5.
In §2 we present our method, in §3 we show our results on GWTC-3, in §4 we present validation tests using simulated GW events, and discuss spin measurements in §5. In §6 we discuss astrophysical implications of our findings, and in §8 we outline our summary and conclusions. Uncertainties throughout this work are at the 90% Credible Interval (CI).
2 Methods
We use hierarchical Bayesian population inference to simultaneously infer the parameters describing the mass and redshift distributions of BBH mergers with parametric models. We provide a brief summary below and refer the reader to Mandel et al. (2019); Thrane & Talbot (2019); Vitale et al. (2021) for an in depth description of the framework.
2.1 Hierarchical Bayesian Inference
@@
$\frac{dN(m_{1},m_{2},z|\Lambda)}{dm_{1}dm_{2}dz}=\frac{dV_{c}}{dz}\bigg{(}\frac{T_{\mathrm{obs}}}{1+z}\bigg{)}\mathcal{R}_{0}p(m_{1},m_{2},z|\Lambda)$ (1)
where $\Lambda$ denotes the hyperparameters describing the population distribution $p(m_{1},m_{2},z|\Lambda)$ of BBH masses and redshifts. Here, $dV_{c}/dz$ is the differential uniform-in-comoving volume element, $T_{\mathrm{obs}}$ is the total observation time, the factor of $1/(1+z)$ converts source-frame time to detector-frame time, and $\mathcal{R}_{0}$ is the local BBH merger rate at $z=0$. The BBH population is modeled through the normalized mass distribution $p(m_{1},m_{2},|\Lambda)$ as well as its redshift evolution $p(z|\Lambda)$, which we parameterize by a power law proportional to $(1+z)^{\kappa}$ as in previous studies (Fishbach et al., 2018; Abbott et al., 2023b).
Given a set of $N_{\mathrm{obs}}$ gravitational wave observations $\{d_{i}\}$, we can calculate the posterior on $\Lambda$ following Ref. Farr et al. (2019) and Mandel et al. (2019):
$\begin{split}p\left(\Lambda|\{d_{i}\}\right)&\propto p(\Lambda)e^{-\mathcal{R}_{0}\beta(\Lambda)}\prod_{i=1}^{N_{\mathrm{obs}}}\Bigg{[}\int\mathcal{L}\left(d_{i}|m_{1},m_{2},z\right)\\ &\times\frac{dN(m_{1},m_{2},z|\Lambda)}{dm_{1}dm_{2}dz}dm_{1}dm_{2}dz\Bigg{]},\end{split}$ (2)
@@
Finally, we marginalize over the local BBH merger rate $\mathcal{R}_{0}$ using a log-Uniform prior (Farr et al., 2019; Abbott et al., 2021a). We also approximate the integral over the individual event likelihoods in Eq. 2 with importance sampling over $N_{i}$ single-event posterior samples generated from single event inference analysis with the default event prior of $\pi(m_{1},m_{2},z,\chi_{\rm{eff}})\propto d_{L}^{2}(1+z)^{2}\frac{dd_{L}}{dz}$. For the $\chi_{\rm{eff}}$ inference in Sec. 5, we compute the default PE prior according to Callister (2021) using the gw-distributions library111publicly available at https://git.ligo.org/reed.essick/gw-distributions.
2.2 Model Comparison
@@ 2.3 Population Models
$p(m_{1},m_{2}|\Lambda)=p(m_{1}|\Lambda)p(m_{2}|\Lambda)f(m_{1},m_{2}|\Lambda).$ (5)
@@
In this work, we consider the following family of models for $p(m|\Lambda)$ as a mixture of power laws $p_{{\rm{pl}}}$ and Gaussian $p_{{\rm{g}}}$ component distributions,
@@
More explicitly, we consider a mixture of two power laws and a single Gaussian component, and therefore we write, {widetext}
@@
3 Results
In this section, we show the inferred population distributions for the masses of BBH mergers using the models we described in Section 2.3. Following the same GWTC-3 (Abbott et al., 2023c) population analysis choices (Abbott et al., 2023b), we only use the BBH events which pass a 1 per year False Alarm Rate (FAR) threshold. Therefore, we analyze a total of 69 confident BBH detections and exclude population outliers such as GW190814 from our analysis. For each event used in our analysis, we make use of the parameter estimation (PE) samples released by LVK (Abbott et al., 2023a). We also make use of the publicly available LVK GWTC-3 sensitivity injection campaign (Abbott et al., 2021b) to estimate the detection fraction $\beta(\Lambda)$ due to selection effects.
@@ 3 Results
First, we note that the PLPL model from a qualitative perspective appears to be a good fit to the GWTC-3 data since the 90% confidence interval on the marginal mass distributions are consistent with Powerlaw+Peak. More quantitatively, we can compute the Bayes factor comparing the PLPL model to the PLB model to obtain an estimate of how well the two models represent the GWTC-3 population with respect to each other. We obtain $\log_{10}\mathcal{B^{\textsc{PLPL}}_{\textsc{PLB}}}=+0.1$, and determine that the PLPL model describes the GWTC-3 observations similarly well to the PLB model, if not marginally better. In other words, the Gaussian bump at around $30-35~{}M_{\odot}$ could be described by a power law which is steeper than that recovered with the PLB model, turning on at a minimum mass of $m_{{\rm{min}},2}\sim 18~{}M_{\odot}$ and maximum mass $m_{{\rm{max}},2}\sim 45-100~{}M_{\odot}$ (at a different location depending on which model is chosen between PLPL and PLPLB), potentially changing the astrophysical interpretation of this population.
@@
4 Validation
We validate our results using a large set of mock GW observations drawn from a known fiducial population model. Specifically, we want to test whether a PLPL or PLPLB model can spuriously fit the data given a BBH population which actually follows a PLB model. We draw simulated BBH events using the default LVK Powerlaw + Peak model (Talbot & Thrane, 2018) for the primary mass distribution. Mass ratios are drawn from a power law with slope $\beta$, and we allow the overall merger rate to evolve with redshift as a power law in $(1+z)^{\kappa}$. Our simulations are constructed to closely follow the inferred GWTC-3 population parameters from Abbott et al. (2023b), hence the fiducial choice for the hyper-parameter values for the simulated population described in Magaña Hernandez & Ray (2024). For reference, the marginalized component mass distributions for our simulated population are shown in Figure 3.
@@ 4 Validation
The goal of our validation study is to understand the distribution of $\log_{10}\mathcal{B}$ obtained on GWTC-3-sized catalogs. To do this, we use 69 mock BBH observations (drawn from our simulated catalog) and repeat this procedure so that we have 100 independent realizations in total. We then perform hierarchical population inference on each of our GWTC-3-sized catalogs with each of the population models considered in this work. We then compute the corresponding $\log_{10}\mathcal{B^{\textsc{PLPL}}_{\textsc{PLB}}}$ and $\log_{10}\mathcal{B^{\textsc{PLPLB}}_{\textsc{PLB}}}$ factors for each catalog realization. In Fig. 4, we show the distribution of $\log_{10}\mathcal{B}$ for both the PLPL and PLPLB models (compared to PLB) and show the corresponding GWTC-3 results from Section 3 for reference.
@@
To quantify the fraction of realizations which can reproduce our results due to clustering of events around $70~{}M_{\odot}$ when no such sub-population is present, in Fig. 5 we show the marginalized posterior distribution on $f_{\rm{g}}$ for both our simulations and GWTC-3. We find that for GWTC-3, $f_{\rm g}>0.01$ is at 95% confidence. For our simulations, only 3% of the realizations have a similar significance of $f_{\rm{g}}>0.01$ at 95% confidence.
Figure 3: Primary (top panel) and secondary (lower panel) component mass distributions (90% confidence interval) from the median mass distribution reconstructions for each of the 100 simulated GWTC-3-sized catalogs described in Section 4 for the PLB (pink), PLPL (blue) and PLB (green) models. For reference, we show the simulated GWTC-3 Powerlaw+Peak like population as the solid black line.
Figure 4: In pink (green) we show the distribution of $\log_{10}\mathcal{B^{\textsc{PLPL}}_{\textsc{PLB}}}$ ($\log_{10}\mathcal{B^{\textsc{PLPLB}}_{\textsc{PLB}}}$) for the simulated GWTC-3-sized catalogs described in Sec. 4. The vertical lines show the corresponding GWTC-3 results.
@@
5 Does spin tell us anything?

### v2

@@ Abstract
We analyze the confident binary black hole (BBH) detections from the third Gravitational-Wave Transient Catalog (GWTC-3) with an alternative mass population model in order to capture features in the mass distribution beyond the Powerlaw + Peak model. We find that the peak of a second power law characterizes the $\sim 30-35~{}M_{\odot}$ bump, such that the data marginally prefers a mixture of two power laws for the mass distribution of binary components over a Powerlaw + Peak model with a Bayes Factor $\log_{10}\mathcal{B}$ of 0.24. This result may imply that the $\sim 30-35~{}M_{\odot}$ feature represents the onset of a second population of BBH mergers (e.g. from a dynamical formation channel) rather than a specific mass feature over a broader distribution. When an additional Gaussian bump is allowed within our power law mixture model, we find a new feature in the BH mass spectrum at $\sim 65-70~{}M_{\odot}$ ($\log_{10}\mathcal{B}$ = 0.29 compared to Powerlaw + Peak). This new feature may be consistent with hierarchical mergers, and constitute $\sim 3\%$ of the BBH population. This model also recovers a maximum mass of $58^{+32}_{-14}~{}M_{\odot}$ for the second power law, consistent with the onset of a pair-instability supernova mass gap.
gravitational-waves, populations
I Introduction
Since the first detection of a binary black hole (BBH) merger in gravitational waves (GW) Abbott et al. (2016a), the field of GW astronomy has entered the era of population studies. During the course of the first three observing runs conducted by LIGO and Virgo, 69 high significance BBH mergers have been detected (Abbott et al., 2023a), and hundreds are expected to be identified throughout the ongoing fourth LIGO-Virgo-KAGRA (LVK) observing run.
Multiple formation channels have been proposed to explain the observed population of compact objects beyond the standard isolated binaries scenario, including formation pathways requiring dynamical interactions in dense star clusters (Rodriguez et al., 2015, 2019), in the disks of Active Galactic Nuclei (AGN) (McKernan et al., 2012; Antonini and Rasio, 2016; McKernan et al., 2018) or in the center of dwarf galaxies (Conselice et al., 2020; Palmese and Conselice, 2021), black holes from metal-free Population III stars (Pop III) (Kinugawa et al., 2014; Hartwig et al., 2016), and primordial black holes (Carr and Hawking, 1974; Clesse and García-Bellido, 2017; Tsai et al., 2021). For in depth reviews, see (Mapelli, 2021; Mandel and Farmer, 2022).
@@ I Introduction
Multiple options for fitting the BBH mass distribution (and more broadly other population properties) have been proposed. The second and third Gravitational-Wave Transient Catalog (GWTC) fiducial analyses include a broken power law or a power law plus peak (Abbott et al., 2021a, b). Recent studies have also modeled the distribution with physics-motivated models (Baxter et al., 2021) or non-parametric approaches (Mandel et al., 2016; Abbott et al., 2021b; Edelman et al., 2023; Ray et al., 2024; Magaña Hernandez and Ray, 2024; Farah et al., 2024a).
Evidence for multiple BH populations includes the presence of a peak in the primary mass distribution at $\sim 30-35~{}M_{\odot}$, which is seemingly too narrow or too unrealistically located to to be explained by various formation channels (Briel et al., 2023; Golomb et al., 2024; Hendriks et al., 2023). Furthermore, BBH from Pop III stars, which may be the seeds of of galaxies’ central black holes still residing in ultra-dwarf galaxies in the nearby Universe, are also predicted to peak around these values (Kinugawa et al., 2014; Conselice et al., 2020; Palmese and Conselice, 2021). Motivated by these theories, we explore whether the $\sim 30-35~{}M_{\odot}$ excess can be explained by a second power law rather than a peak over a broader power law by using the confidently detected BBH observations from the GWTC-3 catalog  (Abbott et al., 2023a)).
Given that we are considering a population model for a formation channel which is possibly of dynamical origin (see Gerosa and Fishbach (2021) for a review), we also explore the presence of an additional high mass hierarchical merger population (Abbott et al., 2020a; Kimball et al., 2020, 2021) that could originate from the merger of two black holes each formed from the merger of another two compact objects in the $30-35~{}M_{\odot}$ peak of the power law.
In §II we present our method, in §III we show our results on GWTC-3, in §IV we present validation tests using simulated GW events, and discuss spin measurements in §V. In §VI we discuss astrophysical implications of our findings, and in §VII we outline our summary and conclusions. Uncertainties throughout this work are at the 90% Credible Interval (CI).
II Methods
We use hierarchical Bayesian population inference to simultaneously infer the parameters describing the mass and redshift distributions of BBH mergers with parametric models. We provide a brief summary below and we refer the reader to review articles for an in depth description of the framework (Mandel et al., 2019; Thrane and Talbot, 2019; Vitale et al., 2021).
II.1 Hierarchical Bayesian Inference
@@
$\frac{dN(m_{1},m_{2},z|\Lambda)}{dm_{1}dm_{2}dz}\propto\frac{dV_{c}}{dz}\bigg{(}\frac{T_{\mathrm{obs}}}{1+z}\bigg{)}\mathcal{R}_{0}(1+z)^{\kappa}p(m_{1},m_{2}|\Lambda)$ (1)
where $\Lambda$ denotes the hyperparameters describing the population distribution $p(m_{1},m_{2}|\Lambda)$ of BBH masses. Here, $dV_{c}/dz$ is the differential uniform-in-comoving volume element, $T_{\mathrm{obs}}$ is the total observation time, the factor of $1/(1+z)$ converts source-frame time to detector-frame time, and $\mathcal{R}_{0}$ is the local BBH merger rate at $z=0$. The BBH population is modeled through the normalized mass distribution $p(m_{1},m_{2},|\Lambda)$ as well as its redshift evolution, which we parameterize by a power law proportional to $(1+z)^{\kappa}$ as in previous studies (Fishbach et al., 2018; Abbott et al., 2023b).
Given a set of $N_{\mathrm{obs}}$ gravitational wave observations $\{d_{i}\}$, we can calculate the posterior on $\Lambda$ following (Farr et al., 2019) and (Mandel et al., 2019):
$\begin{split}p\left(\Lambda|\{d_{i}\}\right)&\propto\frac{p(\Lambda)}{\beta(\Lambda)^{N_{\mathrm{obs}}}}\prod_{i=1}^{N_{\mathrm{obs}}}\Bigg{[}\int\mathcal{L}\left(d_{i}|m_{1},m_{2},z\right)\\ &\times\frac{dN(m_{1},m_{2},z|\Lambda)}{dm_{1}dm_{2}dz}dm_{1}dm_{2}dz\Bigg{]},\end{split}$ (2)
@@
Finally, we marginalize over the local BBH merger rate $\mathcal{R}_{0}$ using a log-Uniform prior (Farr et al., 2019; Abbott et al., 2021a). We also approximate the integral over the individual event likelihoods in Eq. 2 with importance sampling over $N_{i}$ single-event posterior samples generated by the LVK from single event parameter estimation (PE) analyses with prior choices corresponding to $\pi(m_{1},m_{2},z)\propto d_{L}^{2}(1+z)^{2}\frac{dd_{L}}{dz}$ Abbott et al. (2023b). In Sec. V, we analyze additional spin information for each merger encoded in their $\chi_{\rm{eff}}$ measurement (a mass-weighted, aligned, effective spin, see Sec. V for its definition). Therefore, we compute the default PE prior $\pi(m_{1},m_{2},z,\chi_{\rm{eff}})$ according to Callister (2021) using the gw-distributions library111publicly available at https://git.ligo.org/reed.essick/gw-distributions.
II.2 Model Comparison
@@ II.3 Population Models
$p(m_{1},m_{2}|\Lambda)\propto p(m_{1}|\Lambda)p(m_{2}|\Lambda)f(m_{1},m_{2}|\Lambda).$ (5)
@@
In this work, we consider the following family of models for $p(m|\Lambda)$ as a mixture of power laws $p_{{\rm{pl}}}$ and Gaussian $p_{{\rm{g}}}$ component probability distributions,
@@
More explicitly, we consider a mixture of two power laws and a single Gaussian component, and therefore we write,
@@
III Results
In this section, we show the inferred population distributions for the masses of BBH mergers using the models we described in Section II.3. Following the same GWTC-3 (Abbott et al., 2023c) population analysis choices (Abbott et al., 2023b), we only use the BBH events which pass a 1 per year False Alarm Rate (FAR) threshold. Therefore, we analyze a total of 69 confident BBH detections and exclude population outliers such as GW190814 (Abbott et al., 2020b) from our analysis. For each event used in our analysis, we make use of the parameter estimation (PE) samples released by LVK (Abbott et al., 2023a). We also make use of the publicly available LVK GWTC-3 sensitivity injection campaign (Abbott et al., 2021b) to estimate the detection fraction $\beta(\Lambda)$ due to selection effects.
@@ III Results
First, we note that the PLPL model from a qualitative perspective appears to be a good fit to the GWTC-3 data since the 90% confidence interval on the marginal mass distributions are consistent with Powerlaw+Peak. More quantitatively, we can compute the Bayes factor comparing the PLPL model to the PLB model to obtain an estimate of how well the two models represent the GWTC-3 population with respect to each other. We obtain $\log_{10}\mathcal{B^{\textsc{PLPL}}_{\textsc{PLB}}}=+0.24$, and determine that the PLPL model describes the GWTC-3 observations similarly well to the PLB model, if not marginally better. In other words, the Gaussian bump at around $30-35~{}M_{\odot}$ could be described by a power law which is steeper than that recovered with the PLB model, turning on at a minimum mass of $m_{{\rm{min}},2}\sim 18~{}M_{\odot}$ and maximum mass $m_{{\rm{max}},2}\sim 45-100~{}M_{\odot}$ (at a different location depending on which model is chosen between PLPL and PLPLB), potentially changing the astrophysical interpretation of this population.
@@
[…]
IV Validation
We validate our results using a large set of mock GW observations drawn from a known fiducial population model. Specifically, we want to test whether a PLPL or PLPLB model can spuriously fit the data given a BBH population which actually follows a PLB model. We draw simulated BBH events using the default LVK Powerlaw + Peak model (Talbot and Thrane, 2018) for the primary mass distribution. Mass ratios are drawn from a power law with slope $\beta$, and we allow the overall merger rate to evolve with redshift as a power law in $(1+z)^{\kappa}$. Our simulations are constructed to closely follow the inferred GWTC-3 population parameters from Abbott et al. (2023b), hence the fiducial choice for the hyper-parameter values for the simulated population described in Magaña Hernandez and Ray (2024). For reference, the marginalized component mass distributions for our simulated population are shown in Figure 4.
@@ IV Validation
The goal of our validation study is to understand the distribution of $\log_{10}\mathcal{B}$ obtained on GWTC-3-sized catalogs. To do this, we use 69 mock BBH observations (drawn from our simulated catalog) and repeat this procedure so that we have 100 independent realizations in total. We then perform hierarchical population inference on each of our GWTC-3-sized catalogs with each of the population models considered in this work. We then compute the corresponding $\log_{10}\mathcal{B^{\textsc{PLPL}}_{\textsc{PLB}}}$ and $\log_{10}\mathcal{B^{\textsc{PLPLB}}_{\textsc{PLB}}}$ factors for each catalog realization. In Fig. 5, we show the distribution of $\log_{10}\mathcal{B}$ for both the PLPL and PLPLB models (compared to PLB) and show the corresponding GWTC-3 results from Section III for reference.
@@
To quantify the fraction of realizations which can reproduce our results due to clustering of events around $70~{}M_{\odot}$ when no such sub-population is present, in Fig. 6 we show the marginalized posterior distribution on $f_{\rm{g}}$ for both our simulations and GWTC-3. We find that for GWTC-3, $f_{\rm g}>0.01$ is at 96% confidence. For our simulations, only 1 of the realizations have a similar significance of $f_{\rm{g}}>0.01$ at 96% confidence.
Figure 4: Primary (top panel) and secondary (lower panel) component mass distributions (90% confidence interval) from the median mass distribution reconstructions for each of the 100 simulated GWTC-3-sized catalogs described in Section IV for the PLB (pink), PLPL (blue) and PLB (green) models. For reference, we show the simulated GWTC-3 Powerlaw+Peak like population as the solid black line.
Figure 5: In pink (green) we show the distribution of $\log_{10}\mathcal{B^{\textsc{PLPL}}_{\textsc{PLB}}}$ ($\log_{10}\mathcal{B^{\textsc{PLPLB}}_{\textsc{PLB}}}$) for the simulated GWTC-3-sized catalogs described in Sec. IV. The vertical lines show the corresponding GWTC-3 results.
@@
V Does spin tell us anything?